# From Pixels to Words – Towards Native Vision-Language Primitives at Scale

### v1

@@ Abstract
The edifice of native Vision-Language Models (VLMs) has emerged as a rising contender to typical modular VLMs, shaped by evolving model architectures and training paradigms. Yet, two lingering clouds cast shadows over its widespread exploration and promotion: (-) What fundamental constraints set native VLMs apart from modular ones, and to what extent can these barriers be overcome? (-) How to make research in native VLMs more accessible and democratized, thereby accelerating progress in the field. In this paper, we clarify these challenges and outline guiding principles for constructing native VLMs. Specifically, one native VLM primitive should: (i) effectively align pixel and word representations within a shared semantic space; (ii) seamlessly integrate the strengths of formerly separate vision and language modules; (iii) inherently embody various cross-modal properties that support unified vision-language encoding, aligning, and reasoning. Hence, we launch NEO, a novel family of native VLMs built from first principles, capable of rivaling top-tier modular counterparts across diverse real-world scenarios. With only 390M image-text examples, NEO efficiently develops visual perception from scratch while mitigating vision-language conflicts inside a dense and monolithic model crafted from our elaborate primitives. We position NEO as a cornerstone for scalable and powerful native VLMs, paired with a rich set of reusable components that foster a cost-effective and extensible ecosystem.
@@ 1 Introduction
(1) Native VLM Primitive. Native VLMs should embody a unified vision–language primitive that simultaneously integrates encoding, alignment, and reasoning across modalities in one single module. Its design should encompass three principles: (i) a Flexible Position Encoding scheme that generalizes effectively to dynamic spatial structures; (ii) a Multi-Head Native Attention (MHNA) that jointly processes visual–textual connectivity; (iii) Native Rotary Position Embeddings (Native-RoPE) with modality-specific frequencies, preserving compatibility with pretrained LLM’s weights while absorbing original VE’s interaction patterns. Guided by these tenets, we supercharge the fundamental LLM layers with a hybrid attention, expanded heads, and targeted RoPE across modalities, synchronously capturing multi-dimensional relationships for fine-grained and comprehensive correspondence.
@@
We launch NEO, an innovative native VLM that reimagines multi-modal integration from first principles. Unlike typical modular designs, NEO rests on unified primitives that natively encode, align, and reason across modalities, forming coherent pixel–word correspondences from the outset. Through streamlined end-to-end training on 390M image–text samples, NEO acquires strong visual perception and rivals leading modular VLMs of comparable scale across diverse benchmarks. Beyond competitive results, NEO offers reusable components that simplify subsequent development and reduce barriers to promoting native exploration. This reveals that next-generation multimodal systems could also originate from architectures that are native, unified, and intrinsically multimodal.
@@ 2.1 Modular Vision-Language Models
Current Vision–Language Models (VLMs) have converged on a modular paradigm, where a pretrained Vision Encoder (VE) is paired with a Large Language Model (LLM) via lightweight adapters, e.g., projection layers Li et al. (2024a; b) or cross-attention mechanisms Alayrac et al. (2022); Dai et al. (2024). This architecture underlies both leading proprietary systems, including Claude Anthropic (2024; 2025), GPT Hurst et al. (2024); Yang et al. (2023); OpenAI (2025), and Gemini Comanici et al. (2025); DeepMind (2025), as well as prominent open-source frameworks such as InternVL Zhu et al. (2025); Wang et al. (2025b), Qwen-VL Wang et al. (2024a); Bai et al. (2025), and Grok xAI (2024; 2025). By harnessing the complementary strengths of vision and language components, modular architectures, adhering to the “ViT-MLP-LLM” pipeline, achieve unprecedented performance across diverse multimodal benchmarks and have emerged as the dominant design principle in the field.
Despite empirical successes, modular VLMs remain constrained by multi-stage training and heterogeneous structures. Extensive post-training interventions are often required to mitigate rigid inductive biases in pretrained VEs Wang et al. (2024a), which limit resolution flexibility, erode fine-grained details, and blunt sensitivity to features across scales. Besides, pretraining semantic biases and capacity trade-offs between VEs and LLMs collectively impede design simplicity, deployment efficiency, and seamless integration of vision and language, underscoring the urgent need for a monolithic backbone.
@@ 2.2 Native Vision-Language Models
Native VLMs embrace early-fusion integration rather than grafting VEs onto LLMs. Early Fuyu Bavishi et al. (2023), EVE Diao et al. (2024), and SOLO Chen et al. (2024d), embed image patches via linear projections, whereas Chameleon Team (2024), MoMA Lin et al. (2024), and MoT Liang et al. (2024) transform images into symbolic sequences via discrete tokenizers. Later studies Luo et al. (2024); Diao et al. (2025); Li et al. (2025b); Luo et al. (2025); Li et al. (2025a) leverage Mixture-of-Experts (MoE) or Divide-and-Conquer (DaC) strategies to suppress vision-language interference, while others Diao et al. (2024); Li et al. (2025b); Wang et al. (2025a); Li et al. (2025a) upgrade visual encoder supervision to accelerate the acquisition of visual concepts. Empirical evidence Beyer et al. (2024); Luo et al. (2024); Lei et al. (2025) reveals that, with sufficient data and progressive training, native VLMs rapidly approach modular counterparts, corroborating recent scaling-law insights Shukor et al. (2025b; a). Besides, recent methods Tao et al. (2025); Yan et al. (2025); Xiao et al. (2025) indicate that multi-modality encoding modules with the LLM or VE style slightly resolve vision-language misalignment, yet fail to fully integrate the distinct properties of each modality.
@@
Notably, NEO redefines native VLMs as a unibody system built from first-principle primitives. Every component—from image patch encoding, attention mechanism to rotary position embeddings—ensures full compatibility with the intrinsic modeling patterns of VEs and LLMs. Meanwhile, NEO evolves modular VLM strengths via the modality-agnostic pre-Buffer and end-to-end training, dramatically enhancing pixel-word alignment and pushing the frontier of native VLM research.
@@ 3.1 Model Architecture
Patch and Word Embeddings. Given an image $\bm{\mathrm{I}}$, we convert it into token sequences via a lightweight Patch Embedding Layer (PEL) with two Convolutional layers (Conv1–2) Krizhevsky et al. (2012) and a Gaussian Error Linear Unit (GELU) Hendrycks & Gimpel (2016). For input text $\bm{\mathrm{T}}$, we encode it into word tokens using the original LLM Tokenizer as Word Embedding Layer (WEL):
@@
where $\bm{x_{v}}\in\mathbb{R}^{(h\times w)\times d}$ / $\bm{x_{t}}\in\mathbb{R}^{n\times d}$ denote visual / textual tokens, and $\bm{\mathrm{PE}}$ is 2D Sinusoidal Positional Encoding Dosovitskiy et al. (2021). The stride of Conv1 / Conv2 is 16 / 2, i.e., each visual token corresponds to a $32\times 32$ image patch. Notably, Conv2 performs token folding like pixel unshuffle Chen et al. (2024e), with the special <img> and </img> tokens inserted at the boundaries of visual tokens, while mapping position and patch embeddings into a unified space. Afterward, visual and textual tokens are merged and propagated through the unified backbone.
@@
Native VLM Primitive. It adopts RMSNorm Zhang & Sennrich (2019) and SwiGLU Dauphin et al. (2017) consistent with the original LLM layers. Unlike prior methods Wang et al. (2024a); Bai et al. (2025); Zhu et al. (2025); Wang et al. (2025b) that collapse visual tokens into 1D representations or merely reallocate pre-trained LLM head dimensions across temporal (T), height (H), and width (W), we expand Query (Q) and Key (K) head dimensions while fully decoupling H, W, and T relations in Figure 3(1), causing an extra 10% parameter counts over the original Transformer block. The original T dimension is preserved, and additional H and W dimensions are added with their respective QK normalization Yang et al. (2025). Crucially, the linear weights of K for H and W channels are zero-initialized, and attention scale matches that of LLMs, maintaining the LLM pre-training paradigm and progressively activating multimodal capabilities for visual spatial relationships.
@@
This philosophy aligns with our Native Rotary Position Embedding (Native-RoPE) in Figure 3(2), which eliminates correlations between H / W and T indexes, while decoupling channel allocations of H, W, and T under the original LLM frequency. (1) Index Allocation: For text, T index is retained while H / W indexes are zeroed. Notably, Native-RoPE is equivalent to 1D-RoPE before training. For images, each visual token has a constant T index, with unique H / W indexes encoding spatial location. Videos, treated as sequences of frames, increment T index per frame, while H / W indexes follow the same spatial scheme as images. In multimodal inputs, each modality’s T index starts from the maximum ID of the preceding modality, ensuring continuous and unambiguous positional encoding across modalities. This serves two purposes: (a) For image pairs, H / W indexes start independently from (0,0), and tokens at corresponding positions share identical dependency, strongly reinforcing correlations and interactions across matching regions Liao et al. (2025); Wu et al. (2025); (b) For image-text pairs, H / W indexes are decoupled from T index and bounded within (0,0) and (H,W), preventing large T index growth from disproportionately affecting H / W indexes Wang et al. (2024a); Bai et al. (2025) and thereby keeping spatial dependencies between long-range text and images.
@@
Another key aspect is (2) Channel and Frequency Allocation. Unlike recent 3D-RoPE methods Bai et al. (2025); Wei et al. (2025); Yuan et al. (2025); Liao et al. (2025), we fully decouple the channel and frequency allocation of H, W, and T, equipped with additional Q/K head dimensions for H and W. This resolves two issues: (a) Zeroing H / W indexes for pure text would disrupt the modeling patterns and linguistic capacity of the LLM if restricted to its original channels. Repairing this disruption requires substantial resources; (b) Even with interleaved or segmented reallocation, H and W are theoretically equivalent but are assigned different frequencies. Meanwhile, the RoPE frequency in LLMs is far lower than that of visual encoders in Figure 3(2). This mismatch limits the modeling of relative distances and local semantics. The problem is exacerbated by the disparity in scales, with temporal ranges spanning up to one million and spatial ranges only a few hundred.
@@
where $\beta$ and $\Theta$ indicate the base and rotation frequency across H, W, and T. Notably, temporal T dimension captures both local and long-range relations, whereas spatial H / W dimensions emphasize local dependencies. This also opens avenues for broader applications, e.g., video understanding Wei et al. (2025), multimodal generation Deng et al. (2025b), and editing Deng et al. (2025a).
Inspired by prior works Lei et al. (2025); Deng et al. (2025b); Li et al. (2025a), we treat one single image as a unified meta-unit for autoregressive modeling. To enable this, we propose Native Multi-Modal Attention with mixed masking in Figure 3(c). Text tokens adhere to standard causal attention, attending only to preceding tokens to maintain autoregressive generation. In contrast, image tokens employ full bidirectional attention, enabling exhaustive interactions among all visual tokens, akin to a visual encoder. This design captures rich spatial and contextual dependencies within images and facilitates vision-language correspondences, thereby supporting complex multimodal reasoning. We use FlexAttention Dong et al. (2024) to minimize memory overhead and increase throughput, as variable-length block-wise attention is fully optimized through CUDA kernel modifications.
Pre-Buffer and Post-LLM. Drawing on modular designs Bai et al. (2025); Wang et al. (2025b), we split NEO into encoding and reasoning components at the outset. In contrast, we build a modality-shared pre-Buffer via native primitives to map vision and language into a unified representation space. We further design a post-LLM via native primitives to absorb the powerful language proficiency and reasoning capabilities of LLMs. This, in turn, promotes deep pixel-word integration within the pre-Buffer—a deliberate design choice to ensure rich multimodal alignment while minimizing disturbance to the LLM. The layer depth in the pre-Buffer and post-LLM primarily refers to the model parameters of existing VEs and LLMs, ensuring a relatively fair comparison while balancing accuracy and efficiency. Crucially, this separation exists only during pre-training to boost visual learning; during mid-training and supervised fine-tuning, the components are upgraded to a monolithic backbone, allowing the VLM to automatically allocate capacity for encoding, alignment, and reasoning.
@@
3.2 Training Procedure
Figure 4 illustrates the whole training pipeline, which proceeds through three progressive stages: pre-training, mid-training, and supervised fine-tuning. The entire model is optimized end-to-end.
Pre-Training Stage. In this phase, NEO acquires fundamental visual concepts and contextual dependencies from scratch, guided by pre-trained patterns from LLMs. Training leverages 345M web-scale and synthetic image-caption pairs, including 100M English and 20M Chinese pairs from LAION-400M Schuhmann et al. (2021), 150M English pairs from COYO-700M Byeon et al. (2022), 20M long-caption examples from BLIP3o Chen et al. (2025), and 5M short-caption pairs from OpenImages Kuznetsova et al. (2018), recaptioned with a pre-trained InternVL2-8B model. The dataset is further enriched with 30M samples from LAION-COCO Schuhmann et al. (2022) and 20M examples from Wukong Gu et al. (2022) with rich Optical Character Recognition (OCR) annotations. A 3:7 ratio of language to multi-modal data is incorporated to reconstruct text projections in the pre-Buffer. Only the patch embedding layer, the pre-Buffer, and additional QK linear weights and normalization, along with H and W, are trainable and optimized with a simple next-token prediction objective. Notably, the new QK heads not only counteract the LLM’s strong language bias that limits visual specialization but also safeguard its capabilities against the effects of low-quality data.
Mid-Training Stage. The objective at this stage is to strengthen the alignment between visual and linguistic capabilities while progressively enhancing recognition of high-resolution images, complex scenes, object scales, spatial grounding, and compact OCR content. The training data is drawn from the pre-training corpus of InternVL-1.5 Chen et al. (2024f), comprising 40M samples across image captioning, conversation, detection, and OCR data, which account for approximately 66%, 11%, 8%, and 15% of the total, respectively. A 3:7 ratio of language to multi-modal data is again applied. The entire architecture is updated with the same loss functions to consolidate vision-language alignment, thereby equipping NEO with the foundational abilities required for various visual scenarios.
Supervised Fine-Tuning Stage. During the SFT stage, NEO’s ability to follow complex linguistic instructions and varied dialogue patterns is further enhanced, a critical step towards real-world deployment. The full network is optimized across diverse high-quality, multi-source instruction datasets. Following Mono-InternVL Luo et al. (2024), we employ about 4M bilingual instructions for supervised learning, covering tasks such as visual question answering, multimodal dialogue, mathematics, and knowledge reasoning. Details of the instruction data are provided in the Appendix.

### v2

@@ Abstract
The edifice of native Vision-Language Models (VLMs) has emerged as a rising contender to typical modular VLMs, shaped by evolving model architectures and training paradigms. Yet, two lingering clouds cast shadows over its widespread exploration and promotion: (-) What fundamental constraints set native VLMs apart from modular ones, and to what extent can these barriers be overcome? (-) How to make research in native VLMs more accessible and democratized, thereby accelerating progress in the field. In this paper, we clarify these challenges and outline guiding principles for constructing native VLMs. Specifically, one native VLM primitive should: (i) effectively align pixel and word representations within a shared semantic space; (ii) seamlessly integrate the strengths of formerly separate vision and language modules; (iii) inherently embody various cross-modal properties that support unified vision-language encoding, aligning, and reasoning. Hence, we launch NEO, a novel family of native VLMs built from first principles, greatly narrowing the gap with top-tier modular counterparts across diverse real-world scenarios. With 390M image-text examples, NEO efficiently develops visual perception from scratch while mitigating vision-language conflicts inside a dense and monolithic model crafted from our elaborate primitives. We position NEO as a cornerstone for scalable and powerful native VLM development, paired with a rich set of reusable components that foster a cost-effective and extensible ecosystem.
@@ 1 Introduction
(1) Native VLM Primitive. Native VLMs should embody a unified vision–language primitive that simultaneously integrates encoding, alignment, and reasoning across modalities in one single module. Its design should encompass three principles: (i) a Flexible Position Encoding scheme that generalizes effectively to dynamic spatial structures; (ii) a Multi-Head Native Attention (MHNA) that jointly processes visual–textual connectivity; (iii) Native Rotary Position Embedding (Native-RoPE) that preserves compatibility with pretrained LLM while absorbing VE’s interaction patterns. Guided by these tenets, we evolve the LLM blocks into native VLM primitives with brand-new RoPE designs and modality-aware interaction patterns, thereby capturing multi-dimensional relationships for fine-grained and comprehensive correspondence from an intrinsically multimodal perspective.
@@
We launch NEO, an innovative native VLM that reimagines multi-modal integration from first principles. Unlike typical modular designs, NEO rests on unified primitives that natively encode, align, and reason across modalities, forming coherent pixel–word correspondences from the outset. Through streamlined end-to-end training on 390M image–text samples, NEO acquires strong visual perception and approaches leading modular VLMs of comparable scale across diverse benchmarks. Beyond these results, NEO offers reusable components that simplify subsequent development and reduce barriers to promoting native exploration. This reveals that next-generation multimodal systems could also originate from architectures that are native, unified, and intrinsically multimodal.
@@ 2.1 Modular Vision-Language Models
Current Vision–Language Models (VLMs) have converged on a modular paradigm, where a pretrained Vision Encoder (VE) is paired with a Large Language Model (LLM) via lightweight adapters, e.g., projection layers (Li et al., 2024a; b) or cross-attention mechanisms (Alayrac et al., 2022; Dai et al., 2024). This encoder-based architecture underlies various popular and leading vision-language systems, including InternVL (Zhu et al., 2025; Wang et al., 2025b), Qwen-VL (Wang et al., 2024a; Bai et al., 2025), Seed-VL (Guo et al., 2025), GLM-V (Hong et al., 2025), and Grok (xAI, 2024; 2025). By harnessing the complementary strengths of vision and language components, modular architectures, adhering to the “ViT-MLP-LLM” pipeline, achieve unprecedented performance across diverse multimodal benchmarks and have emerged as the dominant design principle in the field.
Despite empirical successes, modular VLMs remain constrained by multi-stage training and heterogeneous structures. Extensive post-training interventions are often required to mitigate rigid inductive biases in pretrained VEs (Wang et al., 2024a), which limit resolution flexibility, erode fine-grained details, and blunt sensitivity to features across scales. Besides, pretraining semantic biases and capacity trade-offs between VEs and LLMs collectively impede design simplicity, deployment efficiency, and seamless integration of vision and language, underscoring the urgent need for a monolithic backbone.
@@ 2.2 Native Vision-Language Models
Native VLMs embrace early-fusion integration rather than grafting VEs onto LLMs. Early Fuyu (Bavishi et al., 2023), EVE (Diao et al., 2024), and SOLO (Chen et al., 2024d), embed image patches via linear projections, whereas Chameleon (Team, 2024), MoMA (Lin et al., 2024), and MoT (Liang et al., 2024) transform images into symbolic sequences via discrete tokenizers. Later studies (Luo et al., 2024; Diao et al., 2025; Li et al., 2025b; Luo et al., 2025; Li et al., 2025a) leverage Mixture-of-Experts (MoE) or Divide-and-Conquer (DaC) strategies to suppress vision-language interference, while others (Diao et al., 2024; Li et al., 2025b; Wang et al., 2025a; Li et al., 2025a) upgrade visual encoder supervision to accelerate the acquisition of visual concepts. Empirical evidence (Beyer et al., 2024; Luo et al., 2024; Lei et al., 2025) reveals that, with sufficient data and progressive training, native VLMs rapidly approach modular counterparts, corroborating recent scaling-law insights (Shukor et al., 2025b; a). Besides, recent methods (Tao et al., 2025; Yan et al., 2025; Xiao et al., 2025) indicate that multi-modality encoding modules with the LLM or VE style slightly resolve vision-language misalignment, yet fail to fully integrate the distinct properties of each modality.
@@
Notably, NEO redefines native VLMs as a unibody system built from first-principle primitives. Every network component, from native rotary position embeddings to multi-modality interaction patterns, ensures full compatibility with the intrinsic modeling patterns of VEs and LLMs. Meanwhile, NEO differs from existing modular VLMs via modality-agnostic pre-Buffer and end-to-end training, dramatically enhancing pixel-word alignment and pushing the frontier of native VLM research.
@@ 3.1 Model Architecture
Patch and Word Embeddings. Given an image $\bm{\mathrm{I}}$, we convert it into token sequences via a lightweight Patch Embedding Layer (PEL) with two Convolutional layers (Conv1–2) (Krizhevsky et al., 2012) and a Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel, 2016). For input text $\bm{\mathrm{T}}$, we encode it into word tokens using the original LLM Tokenizer as Word Embedding Layer (WEL):
@@
where $\bm{x_{v}}\in\mathbb{R}^{(h\times w)\times d}$ / $\bm{x_{t}}\in\mathbb{R}^{n\times d}$ denote visual / textual tokens, and $\bm{\mathrm{PE}}$ is 2D Sinusoidal Positional Encoding (Dosovitskiy et al., 2021). The stride of Conv1 / Conv2 is 16 / 2, i.e., each visual token corresponds to a $32\times 32$ image patch. Notably, Conv2 performs token folding like pixel unshuffle (Chen et al., 2024e), with the special <img> and </img> tokens inserted at the boundaries of visual tokens, while mapping position and patch embeddings into a unified space. Afterward, visual and textual tokens are merged and propagated through the unified backbone.
@@
Native VLM Primitive. It adopts RMSNorm (Zhang and Sennrich, 2019) and SwiGLU (Dauphin et al., 2017) consistent with LLM layers. Unlike prior methods that collapse visual tokens into 1D representations (Zhu et al., 2025; Wang et al., 2025b) or merely reallocate pre-trained LLM head dimensions across temporal (T), height (H), and width (W) (Wang et al., 2024a; Bai et al., 2025), we enlarge Query (Q) and Key (K) head dimensions and decouple H, W, and T relations in Figure 3(1), adding $\sim$10% more parameters over the raw Transformer block. The T dimension is retained, and new H and W dimensions are added with their respective QK normalization (Yang et al., 2025).
@@
This philosophy aligns with our Native Rotary Position Embedding (Native-RoPE) in Figure 3(2). (a) Index Allocation: For text, T index is retained while H / W indexes are zeroed. For images, each visual token has a constant T index, with unique H / W indexes encoding spatial location. Videos, treated as sequences of frames, increment T index per frame, while H / W indexes follow the same spatial scheme as images. In multimodal inputs, each modality’s T index starts from the maximum ID of the preceding modality, ensuring continuous and unambiguous positional encoding across modalities. This serves two purposes: (-) For image pairs, H / W indexes start independently from (0,0), and tokens at corresponding positions share identical dependency, strongly reinforcing correlations and interactions across matching regions (Liao et al., 2025; Wu et al., 2025); (-) For image-text pairs, H / W indexes are decoupled from T index and bounded within (0,0) and (H,W), preventing large T index growth from disproportionately affecting H / W indexes (Wang et al., 2024a; Bai et al., 2025) and thereby keeping spatial dependencies between long-range text and images.
@@
Another key aspect is (b) Channel and (c) Frequency Allocation. Unlike recent 3D-RoPE methods (Bai et al., 2025; Wei et al., 2025; Yuan et al., 2025; Liao et al., 2025), we fully decompose the channel and frequency allocation of H, W, and T, equipped with additional Q/K head dimensions for H and W. This resolves two issues: (-) Zeroing H / W indexes for pure text would disrupt the modeling patterns and linguistic capacity of the LLM if restricted to its original channels. Repairing this disruption requires substantial resources; (-) Even with interleaved or segmented reallocation, H and W are theoretically equivalent but are assigned different frequencies. Meanwhile, the RoPE frequency in LLMs is far lower than that of visual encoders in Figure 3(2c). This mismatch limits the modeling of relative distances and local semantics. The problem is exacerbated by the disparity in scales, with temporal ranges spanning up to one million and spatial ranges only a few hundred.
@@
where $\beta$ and $\Theta$ indicate the base and rotation frequency across H, W, and T. Notably, temporal T dimension captures both local and long-range relations, whereas spatial H / W dimensions emphasize local dependencies. This also opens avenues for broader applications, e.g., video understanding (Wei et al., 2025), multimodal generation (Deng et al., 2025b), and editing (Deng et al., 2025a).
Inspired by prior works (Lei et al., 2025; Deng et al., 2025b; Li et al., 2025a; Beyer et al., 2024), we also treat one single image as a unified meta-unit for autoregressive modeling, denoted as Native Multi-Modal Attention with mixed masking in Figure 3(3). Text tokens adhere to standard causal attention, attending only to preceding tokens to maintain autoregressive generation. In contrast, image tokens employ full bidirectional attention, enabling exhaustive interactions among all visual tokens, akin to a visual encoder. This design captures rich spatial and contextual dependencies within images and facilitates vision-language correspondences, thereby supporting complex multimodal reasoning. We use FlexAttention (Dong et al., 2024) to minimize memory overhead and increase throughput, as variable-length block-wise attention is fully optimized through CUDA kernel modifications.
Pre-Buffer and Post-LLM. In Figure 3(4), we develop a modality-shared pre-Buffer to translate pixel–word inputs into a unified representation with minimal disturbance to the post-LLM, which inherits the linguistic proficiency and reasoning capabilities of pre-trained LLM. The layer depths $L_{1}$ and $L_{2}$ primarily refer to parameter counts and scaling properties (Tian et al., 2025) of existing VEs and LLMs to balance accuracy and efficiency. Here, we formulate one primitive $\Phi^{l}$ as follows:
$\bm{x_{m}^{l^{\prime}}}=\bm{x_{m}^{l}}+\text{MHNA}(\text{RMSNorm}(\bm{x_{m}^{l}})),\quad\bm{x_{m}^{l+1}}=\bm{x_{m}^{l^{\prime}}}+\text{FFN}(\text{RMSNorm}(\bm{x_{m}^{l^{\prime}}})),$ (3)
where $\bm{m}\in\{\bm{v},\bm{t}\}$ indicates input modality. Besides, $\{\Phi^{0},...,\Phi^{L_{1}-1}\}$ and $\{\Phi^{L_{1}},...,\Phi^{L_{1}+L_{2}-1}\}$ denotes pre-Buffer and post-LLM, respectively. Notably, we randomly initialize the entire pre-Buffer, while the post-LLM inherits RMSNorm, Feed-Forward Network (FFN), and Q/K/QK-Norm parameters along the temporal dimension from a pretrained LLM. The temporal Q is reused to initialize Q for the H and W dimensions, their K weights are zero-initialized, and the corresponding QK-Norm is initialized with $\beta=0$ and $\gamma=1$. We further match the attention scaling to that of the pretrained LLM, thereby preserving its pre-training paradigm from the outset and enabling a progressive emergence of multimodal spatial reasoning within the post-LLM. Crucially, this separation exists only during pre-training. After that, these components are merged into a monolithic backbone that autonomously allocates capacity for encoding, alignment, and reasoning.
3.2 Training Procedure
Figure 4 illustrates the whole training pipeline, where the entire model is optimized end-to-end.
@@ 3.2 Training Procedure
Pre-Training Stage. In this phase, NEO acquires fundamental visual concepts and contextual dependencies from scratch, guided by pre-trained patterns from LLMs. Training leverages 345M web-scale and synthetic image-caption pairs, including 100M English and 20M Chinese pairs from LAION-400M (Schuhmann et al., 2021), 150M English pairs from COYO-700M (Byeon et al., 2022), 20M long-caption examples from BLIP3o (Chen et al., 2025), and 5M short-caption pairs from OpenImages (Kuznetsova et al., 2018), recaptioned with a pre-trained InternVL2-8B model. The dataset is further enriched with 30M samples from LAION-COCO (Schuhmann et al., 2022) and 20M examples from Wukong (Gu et al., 2022) with rich Optical Character Recognition (OCR) annotations. A 3:7 ratio of language to multi-modal data is incorporated to reconstruct text projections in the pre-Buffer. Only the patch embedding layer, the pre-Buffer, and extra QK linear weights and normalization in post-LLM, along with H and W, are optimized with a simple next-token prediction objective. Notably, the new QK heads not only counteract the LLM’s strong language bias that limits visual specialization but also safeguard its capabilities against the effects of low-quality data.
Mid-Training Stage. The objective at this stage is to strengthen the alignment between visual and linguistic capabilities while progressively enhancing recognition of high-resolution images, complex scenes, object scales, spatial grounding, and compact OCR content. The training data is drawn from the pre-training corpus of InternVL-1.5 (Chen et al., 2024f), comprising 40M samples across image captioning, conversation, detection, and OCR data, which account for approximately 66%, 11%, 8%, and 15% of the total, respectively. A 3:7 ratio of language to multi-modal data is again applied. The entire architecture is updated with the same loss functions to consolidate vision-language alignment, thereby equipping NEO with the foundational abilities required for various visual scenarios.
Supervised Fine-Tuning Stage. During the SFT stage, NEO’s ability to follow complex linguistic instructions and varied dialogue patterns is further enhanced, a critical step towards real-world deployment. The full network is optimized across diverse high-quality, multi-source instruction datasets. Following Mono-InternVL (Luo et al., 2024), we employ about 4M bilingual instructions for supervised learning, covering tasks such as visual question answering, multimodal dialogue, mathematics, and knowledge reasoning. Details of the instruction data are provided in the Appendix.